Circular No. 8587

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 P3 (SWAN) Several people reported the appearance of a possible comet in ultra-

violet SWAN website images taken with the SOHO spacecraft, including H. Sato (Sukagawa, Fukushima-ken, Japan; via S. Nakano, Sumoto, Japan), M. Suzuki (Utsunomiya, Tochigi, Japan), M. Mattiazzo (Adelaide, S. Australia), M. Jäger (Vienna, Austria), and V. Bezugly (Dnepropetrovsk, Ukraine). Following requests for confirmation by the Central Bureau, A. Hale (Cloudcroft, NM) reported a visual observation of the object on Aug. 25.142 UT with a 0.41-m reflector, which had a moderately condensed 3′ coma and a guessed total mag of roughly 9.5; his observation was combined with the necessarily very rough SWAN positions below measured by Sato to produce an ephemeris for the ‘NEO Confirmation Page’. Available positions follow (Aug. 4 and 7 positions from Mattiazzo):

2005 UT  α_{2000}  δ_{2000}  Mag. Observer
Aug. 4  11^{h}12.^{m}2  +15^{o}20′  SWAN
7  11 16.1  +18 31  _{”}
9  11 20.9  +21 08  _{”}
14  11 30.3  +28 07  _{”}
16  11 31.8  +30 11  _{”}
18  11 33.3  +32 34  _{”}
21  11 33.7  +35 22  _{”}
24.83145 11 33 54.^{s}79 +37 38 23′′.5 11.8 Sanchez
24.84571 11 33 53.86 +37 38 57.1  _{”}
24.87452 11 33 53.77 +37 40 21.0  _{”}
25.142 11 33 50 +37 52.0 Hale
26.12613 11 33 38.53 +38 40 13.6 14.4 McGaha
26.12675 11 33 38.51 +38 40 14.8  _{”}
26.43616 11 33 33.75 +38 54 37.7  Kadota
26.43990 11 33 33.73 +38 54 45.8  _{”}
26.44168 11 33 33.67 +38 54 51.4 11.0  _{”}

S. Sanchez, R. Stoss, J. Nomen (Mallorca). 0.07-m f/2.8 refractor + CCD.

Clearly diffuse. J. E. McGaha (Tucson). 0.36-m f/10.0 Schmidt-Cassegrain reflector +

CCD. Inner coma of size 15′′ × 10′′, outer coma of size 45′′ × 32′′ (elongated in both cases in p.a. 225^{o}.

K. Kadota (Ageo, Japan). 0.25-m f/5.0 reflector + CCD. Diffuse coma of diameter 2′ with central condensation, no tail. Comm. by Nakano.

2005 August 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green